OpenStudy (anonymous):

slove and show work please.....2(x+1)/4=(3x-2)...i know i have to find the lcd which i think is 12.....but then i get stuck.

OpenStudy (agreene):

is the 2 in 2(x+1)/4 also being divided by 4? that is: [2(x+1)]/4 ?

OpenStudy (anonymous):

yes

OpenStudy (agreene):

[2(x+1)]/4 = 3x-2 Multiply both sides by 4 2(x+1) = 6x - 6 Distribute the 2 on the left side 2x+2 = 6x -6 Add six to both sides, and subtract 2x 8 = 4x Divide both sides by 4 2 = x :) Ask questions.

OpenStudy (agreene):

I cant multiply apparently... same steps but yeah 2(x+1) = 6x - 8 2x+2=6x-8 add 8 to both sides and sub 2x -6=4x divide -3/2 = x

OpenStudy (anonymous):

the answer in the back of the book is 1.......

OpenStudy (agreene):

Wow... again, I'm bad at multiplying, let me start over, lol

OpenStudy (agreene):

[2(x+1)]/4 = 3x-2 Multiply both sides by 4 2(x+1) = 12x - 8 Distribute the 2 on the left 2x+2 = 12x - 8 Add 8 and subtract 2x from both sides 10 = 10 x divide both sides by 10 1 = x Your book is right, I'm being dumb this morning, sorry.

OpenStudy (anonymous):

thanks for your help....